Calculate the molarity of 32.25 grams of NaCl in 2.0 Liters of solution.
1
Expert's answer
2020-04-20T15:07:53-0400

"\frac{m}{M}=\frac{32.25}{58.443}=0.55 mole"

"molarity=\frac{n}{V}=\frac{0.55}{2.0}=0.275mole/L"